OpenStudy (anonymous):

Linear approximation tan(44)

hartnn (hartnn):

tan(45-1)

hartnn (hartnn):

f(x) = tan x

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

f'(x)=sec^2(x)

OpenStudy (anonymous):

f(a)-f'(a)(x-a)

OpenStudy (anonymous):

tan45=1

OpenStudy (anonymous):

sec^2(45)=2

hartnn (hartnn):

yup, you are going good.

OpenStudy (anonymous):

\[1+2(x-a)\]

OpenStudy (anonymous):

this is where im stuck

OpenStudy (anonymous):

a=45?

OpenStudy (anonymous):

x=44?

hartnn (hartnn):

you took a= 45, thats correct, take x= 1

OpenStudy (anonymous):

-1 degrees?

hartnn (hartnn):

what is in the left side of f(a)-f'(a)(x-a) ??

OpenStudy (anonymous):

f(a) = tan45

OpenStudy (anonymous):

f'(a)=sec^2(45)

OpenStudy (anonymous):

x=44 degrees?

hartnn (hartnn):

i meant its f(x-a) = f(a)-f'(a)(x-a) right? a=45 is correct

hartnn (hartnn):

take x=1

OpenStudy (anonymous):

ok so, 1-45?

OpenStudy (anonymous):

-44 degrees?

hartnn (hartnn):

and tan(-y)=-tan y

hartnn (hartnn):

its actually f(x)= f(a) +f'(a) (x-a)

OpenStudy (anonymous):

i hate the x-a part

OpenStudy (turingtest):

the statement is true for x in the *neighborhood* of a remember that f(a) is the value that is easy to find, and f(x) is the value that we are going to find at the end

hartnn (hartnn):

f(a) + f'(a)(x-a) ......a=45, x=44 =1 + 2*tan(44 - 45)

OpenStudy (anonymous):

tan(-1degrees)

hartnn (hartnn):

yup, its -0.017

hartnn (hartnn):

==1 + 2*tan(44 - 45) =1 + 2( -0.017) = 1 - 0.034 =?

OpenStudy (anonymous):

.968

OpenStudy (anonymous):

if the book has .965, thats close enough?

hartnn (hartnn):

1-0.034 = 0.966 this is close enough

OpenStudy (anonymous):

also, am I supposed to use a calc to find tan-1

hartnn (hartnn):

i don't see any other way out. for more accuracy, tan(-1)=-0.174 then tan 44=0.965

OpenStudy (anonymous):

ahhh, ok, that clears it all up for me, i just need to remember to place tan(x-a) thanks @hartnn !!!!!!

hartnn (hartnn):

welcome ^_^

OpenStudy (goformit100):

(^^^)

OpenStudy (dean.shyy):

Take a look at this: http://is.gd/LCUwSA